A machine gun has a mass of 200 Kg. It fires 35 gm bullets at the rate of 40 bullets per second with a speed of 400 m/s.The force needed to keep the gun fixed is_____________________

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sindhuja

conser the liner momentum

200*0 + 35/1000 *0  =200 *v  +35/1000  *400

v=-57.142 m/sec

we know that F=rate of change of linear momentum

                     =(0-(-200*57.142))/(1/40)

F= 457136 N

  =45.7kN


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin